TWO POSITIVE POINT CHARGES PLACED AT A DISTANCE ‘b’ APART HAVE SUM Q.FOR WHAT VALUES OF THE CHARGES IS THE COULOMB FORCE BETWEEN THEM MAXIMUM?

Grade:Select Grade

2 Answers

siddharth gupta
9 years ago
METHOD 1:
LET THE CHARGES ON THE TWO CHARGED PARTICLES BE Q1 AND Q2 
ELECTROSTATIC FORCE BETWEEN THEM=K(Q1)(Q2)/b2
USING A.M-G.M INEQUALITY FOR Q1Q2 TO BE MAXIMUM , Q1=Q2
i.e Q1=Q2=Q/2.
METHOD 2:
LET THE CHARGES BE Q-X AND X.
F=K(x)(Q-x)/b2
find dF/dx and put it equal to 0 i.eK/b2(Q-2x)=0
i.e Q=2x i.e Q/2=x.
WITH REGARDS 
SIDDHARTH GUPTA
 
Bharat Makkar
9 years ago
THANKS!